Prepare 700 mL of a dextrose solution which 1/3 of it is dextrose and 2/3 is distilled water.

How many mL of water will be used?

How many mL of dextrose will be used?
1
Expert's answer
2017-01-25T08:59:08-0500
We prepared 700 mL of a dextrose solution. Hence, 1/3 of it is dextrose, thus we will have 700 · 1/3 = 700/3 mL of dextrose and 2/3 of it is water, thus we will have 700 · 2/3 = 1400/3 mL of water.

Answer. 700/3 mL of dextrose, 1400/3 mL of water.